# Proxy-Free GFlowNet

### v1

@@ 3.2 Reward-guided pruning and prioritized backward sampling
Although the learned edge reward provides a meaningful signal at the transition level, directly incorporating it into the training objective introduces sensitivity to function approximation errors, which can distort the optimization process and degrade policy performance—mirroring the limitations commonly observed with proxy reward models. To enhance both the robustness and efficiency of the algorithm, we propose two indirect mechanisms for shaping the GFlowNet policy: reward-guided pruning and prioritized backward sampling.
Guided by the learned edge reward, TD-GFN strategically removes low-utility edges from the environment DAG $G=(V,E)$. Specifically, we approximate the distribution of edge rewards using a batch $\mathcal{D}_{R_{E}}=\{R_{E}(s_{i},\text{selected}(\pi_{\psi}(\cdot\mid s_{i})))\}_{i=1}^{\mathcal{B}}$, where $\pi_{\psi}$ denotes the imitation policy obtained during edge reward learning (Algorithm 2), serving as a surrogate for expert behavior. An edge $(s\rightarrow s^{\prime})$ is pruned from the graph if it lies in the low-density region of the reward distribution:
@@
where $K$ is a pruning threshold hyperparameter, and $\text{mean}(\cdot)$, $\text{std}(\cdot)$ denote the empirical mean and standard deviation over the sampled batch $\mathcal{D}_{R_{E}}$.
After applying this criterion, we remove all edges that are disconnected from the root node $s_{0}$, resulting in a pruned subgraph $G^{\prime}=(V^{\prime},E^{\prime})$ and an updated set of terminal nodes $\mathcal{X}^{\prime}$. Notably, pruning is conducted over the full environment DAG, rather than being limited to dataset-observed transitions, enabling the retention of transitions that, while unobserved in the dataset, may potentially lead to high-reward regions.
This pruning approach concentrates learning on informative areas of the state space and improves training efficiency by reducing the complexity of model fitting [30, 31]. The resulting subgraph defines a more compact yet expressive action space that prioritizes structurally meaningful and high-reward trajectories, thereby supporting policy generalization beyond observed data while remaining anchored in reliable behavioral signals.
To train the policy more efficiently on the pruned graph, we additionally propose a prioritized backward sampling mechanism that selects training trajectories from high-value terminal nodes. Terminal states are sampled from the intersection of the pruned terminal set $\mathcal{X}^{\prime}$ and dataset-supported objects $\{s_{T_{i}}\}_{i=1}^{M}$, with probability proportional to their known rewards. From each sampled terminal node $x$, we recursively perform backward sampling toward the root $s_{0}$ using a learned backward policy $\mathcal{P}_{B}$, defined as:
@@
This formulation is consistent with the reward shaping principles in Equation 1. By sharpening policy updates toward the most informative regions of the state space, this prioritized strategy reinforces the GFlowNet inductive bias of allocating sampling effort in proportion to reward, significantly improving sample efficiency [32, 33]. Moreover, when combined with the pruning method described above, it helps mitigate the under-exploitation of terminal nodes discussed in [34].
Finally, we optimize the policy on the pruned DAG using the Flow-Matching (FM) objective [1]. By setting $R(s)=0$ for all non-terminal states $s$, the objective for each sampled trajectory $\tau$ is given by:
@@ 4.1 Hypergrid
We begin with experiments on the Hypergrid task [1], a $D$-dimensional grid environment with side length $H$, resulting in a discrete space of size $H^{D}$. High-reward modes are narrowly concentrated near the $2^{D}$ corners, making the task particularly challenging for both exploration-based and offline learning algorithms. Our main experiments are conducted on a $8^{4}$ Hypergrid ($H=8$, $D=4$) following the setup in [1]. Additionally, in Appendix D.2, we visualize a $8\times 8$ Hypergrid instance and illustrate the TD-GFN workflow on this task.
@@
Due to the lack of existing proxy-free training methods for GFlowNets, we compare TD-GFN against a broad range of baselines trained on offline datasets. These include popular offline RL algorithms such as CQL [35] and IQL [36], imitation learning methods such as Behavior Cloning (BC) and GAIL [28], and the offline GFlowNet variant COFlowNet [22]. We additionally include a baseline termed Dataset-GFN, which trains a GFlowNet directly on offline trajectories without modification, serving as a vanilla proxy-free variant. For both offline RL and imitation learning methods, we adopt their soft RL formulations to enable stochastic policy sampling rather than pure reward maximization. To ensure a fair comparison, we train GAIL using the rebalanced dataset described in Section 3.1, which improves its performance relative to training on the original dataset. Accordingly, we evaluate GAIL using the imitation policy learned via Algorithm 2.
@@
Each algorithm is evaluated using two standard metrics [1, 22, 37, 38]. The first is Empirical L1 Error, which measures the $\mathcal{L}_{1}$ distance between the empirical distribution $\pi(x)$ from model samples and the normalized reward distribution $p(x)=R(x)/Z$, computed as $\mathbb{E}[|\pi(x)-p(x)|]$. The second is Modes Found, defined as the number of distinct reward modes discovered during training.
As shown in Figure 3(a), when trained on an Expert dataset of $1,500$ trajectories collected by a well-trained GFlowNet policy, all baseline methods require at least $30,000$ state visits to uncover all $16$ modes, whereas TD-GFN discovers all modes with fewer than $500$ visits. Although both TD-GFN and COFlowNet achieve the lowest Empirical L1 Error, TD-GFN converges in less than half the number of state visits required by COFlowNet, demonstrating its superior training efficiency. Additional baseline analysis is provided in Appendix D.6.
@@
Figure 3: Performance comparison on the $8^{4}$ Hypergrid task. The top row shows results using $1,500$ trajectories from the Expert and Mixed datasets for policy training, while the bottom row shows results using only $150$ trajectories. The solid line and shaded region represent the mean and standard deviation, respectively.
Robustness to Dataset Uncertainty
@@
We first assess the robustness of TD-GFN under dataset uncertainty by introducing greater variability in the training data. In the Mixed setting, we augment the Expert dataset with an equal number of trajectories generated by a random policy similar to [27, 39], thereby introducing noise. In another setting, we significantly reduce the number of available trajectories—using only one-tenth of the original dataset—to simulate data scarcity as an additional source of uncertainty [40].
@@
Robustness to Behavior Policy Quality
@@
Figure 4: Performance comparison on the $8^{4}$ Hypergrid task. Policies are trained using the Median and Bad datasets, each consisting of $1,500$ trajectories. The solid line and shaded region represent the mean and standard deviation, respectively.
@@ 4.2 Biosequence Design
We next evaluate the performance of TD-GFN on a more challenging task—designing anti-microbial peptides (AMPs), i.e., short protein sequences—introduced in [3]. The experimental setup leverages two datasets curated from the DBAASP database [41]: one for training a proxy model to approximate reward values, and another for training an oracle that provides ground-truth reward labels.
Following the methodology of [1], we train a Proxy-GFN using the proxy model and compare its performance with TD-GFN, as well as with other baseline methods previously evaluated on the Hypergrid task (Section 4.1). All methods are trained on the same dataset used for proxy learning, which includes $3,219$ positive AMPs and $4,761$ non-AMPs.
Figure 5: Comparison of reward and diversity among the top $100$ sequences ranked by reward. The Diversity metric is defined in Appendix C.2. All results are averaged over three random seeds, and error bars indicate the standard deviation.
@@ 4.3 Molecule Design
We further evaluate the effectiveness of our method on a real-world molecular design task introduced by [1]. A training dataset is constructed to simulate historical data produced by human experts, which comprises $1,500$ trajectories generated by a well-trained GFlowNet policy, with corresponding reward values provided by an oracle [1].
Table 1: Comparison of methods on the Molecule Design task. The Convergence column indicates the number of trajectories required for policy convergence. Results are reported as mean $\pm$ standard deviation over three random seeds. Bold numbers indicate the best performance in each column.
@@
CQL $7.069$ $6.643$ $5.401$ 0.803 ×$\bm{\times 10^{4}}$
@@
TD-GFN (Ours) 7.733 $\pm$ 0.036 7.450 $\pm$ 0.037 6.810 $\pm$ 0.035 2.749 ×$\bm{\times 10^{4}}$
Following the approach used in the Biosequence Design task (Section 4.2), we train a proxy model on this dataset using the architecture and methodology commonly adopted in prior GFlowNet research [1, 9, 24, 42, 43], resulting in a baseline referred to as Proxy-GFN. Moreover, [22] proposed an enhanced COFlowNet incorporating quantile matching [38] to improve candidate diversity. Accordingly, we evaluate both the original and quantile-augmented variants (FM/QM).
We report the average top-$k$ reward obtained from $5,000$ samples generated by policies trained with different algorithms, along with the number of training trajectories required for policy convergence, in Table 1. Due to the instability of offline RL methods under sparse rewards, results for CQL and IQL are reported using only their best-performing random seed.
As shown in the table, CQL converges fastest in terms of sample efficiency but struggles to generate a diverse set of high-reward candidates, resulting in substantially lower top-$k$ average rewards than GFN-based and even imitation learning methods. Excluding CQL, TD-GFN achieves the best performance in discovering high-reward molecules with the fewest trajectories, highlighting its training effectiveness and efficiency.
Figure 6: Number of high-reward modes (reward $>7$ and $>7.5$) discovered during training in the Molecule Design task. The solid line and shaded region represent the mean and standard deviation, respectively.
In contrast, although Proxy-GFN performs competitively and is comparable to QM-COFlowNet in terms of reward quality, it requires significantly more trajectory rollouts during training. This further underscores the efficiency of learning directly from offline data. Additional comparisons of training time and computational resource usage are provided in Appendix D.3.
To evaluate the diversity of molecules generated by TD-GFN, we measure the number of distinct high-reward modes identified by each algorithm. Modes are defined as clusters of high-reward molecules based on Tanimoto similarity [44]. As shown in Figure 6, TD-GFN consistently discovers $1.5$ to $2$ times more high-reward modes than the strongest baselines, indicating superior exploration. Notably, despite incorporating DAG pruning, TD-GFN does not overfit to narrow regions of the search space. Instead, it guides the policy toward a broader set of high-reward regions, balancing targeted exploitation with diverse exploration.
@@ 5 Conclusion
We introduced Trajectory-Distilled GFlowNet (TD-GFN), a proxy-free training framework for learning GFlowNets from offline datasets without requiring out-of-distribution reward queries. Leveraging edge-level supervision learned via IRL, TD-GFN consistently guides the policy toward improved performance and training efficiency through DAG pruning and prioritized backward trajectory sampling. Extensive experiments validate the effectiveness of the proposed approach.
@@ Appendix A Algorithm Pseudocode
1:Input: Offline dataset $\mathcal{D}=\{\tau_{i}=(s_{0},\dots,s_{T_{i}},R(s_{T_{i}}))\}_{i=1}^{M}$; number of training iterations $N$; batch size $B$; GFlowNet learning rate $\eta$.
@@
6:Prune the environment DAG $G=(V,E)$ based on the thresholding rule defined in Equation 6.
@@
9:Initialize the GFlowNet flow function $F_{\theta}$.
10:for $n=1$ to $N$ do
@@
16:              Sample $s_{pre}\sim\mathcal{P}_{B}(\cdot|s_{now})$, where $\mathcal{P}_{B}$ is defined in Equation 7 over $G^{\prime}$.
@@
19:         end while // Backward sampling.
@@
21:     end for // Backward sample trajectories.
22:     Optimize $F_{\theta}$ with $\mathcal{T}$ using the FM loss over $G^{\prime}$, as formulated in Equation 8. // Update flow function
@@
24:Extract the GFlowNet policy:
$\mathcal{P}_{F}(s^{\prime}|s)=\pi_{\theta}(s^{\prime}|s)=\begin{cases}\frac{F_{\theta}(s,s^{\prime})}{\sum_{(s\rightarrow s^{\prime\prime})\in E^{\prime}}F_{\theta}(s,s^{\prime\prime})},&\text{if }(s\rightarrow s^{\prime})\in E^{\prime},\\ 0,&\text{otherwise}.\end{cases}$
// Obtain policy from flow
25:Output: Trained GFlowNet policy $\mathcal{P}_{F}$.
@@
6:         Sample a trajectory $\tau_{j}\in\mathcal{D}$ with probability $P(\tau_{j})\propto R(s_{T_{j}})$
@@
8:     end for //Sample from the rebalanced dataset
@@
10:     Compute $\nabla_{\phi}L$ on $\overline{E}$; update $\phi\leftarrow\phi-\beta\nabla_{\phi}L$. // Update discriminator
11:     Compute $\nabla_{\psi}L$ on $\overline{E}$; update $\psi\leftarrow\psi-\alpha\nabla_{\psi}L$ or apply a policy gradient method. // Update policy
@@
13:Extract the edge reward function $R_{E}$ according to Equation 5.
@@
Appendix B Related works

### v2

@@ 3.2 Reward-guided pruning and prioritized backward sampling
Although the learned edge reward provides a meaningful signal at the transition level, using it to directly shape the policy’s gradients introduces sensitivity to function approximation errors—mirroring the error propagation commonly observed with proxy reward models. To enhance both the robustness and efficiency of the algorithm, we propose two novel mechanisms for shaping the GFlowNet policy as illustrated in Figure 3: we perform reward-guided pruning on the environment DAG and subsequently train the GFlowNet using trajectories obtained via prioritized backward sampling on the pruned graph.
Guided by the learned edge reward, TD-GFN removes low-utility edges from the environment DAG $G=(V,E)$. Specifically, we approximate the distribution of edge rewards using a batch $\mathcal{D}_{R_{E}}=\{R_{E}(s_{i},\text{selected}(\pi_{\psi}(\cdot\mid s_{i})))\}_{i=1}^{\mathcal{B}}$, where $\pi_{\psi}$ denotes the imitation policy obtained during edge reward learning (Algorithm 2), serving as a surrogate for expert behavior. An edge $(s\rightarrow s^{\prime})$ is pruned from the graph if it lies in the low-density region of the reward distribution:
@@
where $K$ is a pruning threshold hyperparameter (a sensitivity analysis is provided in Appendix D.4), and $\text{mean}(\cdot)$, $\text{std}(\cdot)$ denote the empirical mean and standard deviation over the sampled batch $\mathcal{D}_{R_{E}}$.
After applying this criterion, we further remove all edges that are disconnected from the root node $s_{0}$, resulting in a pruned subgraph $G^{\prime}=(V^{\prime},E^{\prime})$ and an updated set of terminal nodes $\mathcal{X}^{\prime}$. Notably, pruning is conducted over the full environment DAG, rather than being limited to dataset-observed transitions, enabling the retention of transitions that, while unobserved in the dataset, may potentially lead to high-reward regions.
This pruning approach concentrates learning on informative areas of the state space and improves training efficiency by reducing the complexity of model fitting (Zahavy et al., 2018; Zhang et al., 2020). The resulting subgraph defines a more compact yet expressive action space that prioritizes structurally meaningful and high-reward trajectories, thereby supporting policy generalization beyond observed data while remaining anchored in reliable behavioral signals, as illustrated in Figure 2(c).
To train the policy more efficiently on the pruned graph, we additionally propose a prioritized backward sampling mechanism that constructs training trajectories backwards from terminal nodes. Terminal states are sampled from the intersection of $\mathcal{X}^{\prime}$ and dataset-supported objects $\{s_{T_{i}}\}_{i=1}^{M}$, with probability proportional to their known rewards. From each sampled terminal node $x$, we recursively perform backward sampling toward the root $s_{0}$ using a learned backward policy $\mathcal{P}_{B}$, defined as:
@@
This formulation is consistent with the reward shaping principles in Equation (1). By sharpening policy updates toward regions prioritized by both terminal and edge rewards, this strategy reinforces the GFlowNet inductive bias of allocating sampling effort in proportion to reward, thereby improving sample efficiency. Moreover, when combined with the pruning method described above, it helps mitigate the under-exploitation of terminal nodes discussed in Jang et al. (2024).
Finally, we optimize the policy on the pruned DAG using the sampled trajectories and their corresponding terminal rewards from the offline dataset. Crucially, the gradient-based updates depend only on these recorded, ground-truth values. This approach not only sidesteps the fitting errors inherent in proxy reward models but also insulates the gradients from any potential inaccuracies in the edge rewards learned during the IRL phase. Notably, the contributions of TD-GFN are orthogonal to the specific training strategies employed after trajectory sampling, as discussed in Appendix D.7.
@@ 4.1 Hypergrid
We begin with experiments on the Hypergrid task (Bengio et al., 2021), a $D$-dimensional grid environment with side length $H$, resulting in a discrete space of size $H^{D}$. High-reward modes are narrowly concentrated near the $2^{D}$ corners, making the task particularly challenging for both exploration-based and offline learning algorithms. Our main experiments are conducted on a $8^{4}$ Hypergrid ($H=8$, $D=4$) following the setup in Bengio et al. (2021). We also provided results on a larger Hypergrid in Appendix D.2. Additionally, in Appendix D.3, we visualize a $8\times 8$ Hypergrid instance and illustrate the TD-GFN workflow on this task.
@@
Due to the nascent nature of proxy-free GFlowNet training, we compare TD-GFN against a broad range of baselines. These include the main existing proxy-free GFlowNet method, COFlowNet (Zhang et al., 2025b), as well as popular offline RL algorithms (soft RL version) such as CQL (Kumar et al., 2020) and IQL (Kostrikov et al., 2021), imitation learning methods such as Behavior Cloning (BC) and GAIL (Ho & Ermon, 2016). We additionally include a baseline termed Dataset-GFN, which trains a GFlowNet directly on offline trajectories without modification, serving as a vanilla proxy-free variant. To ensure a fair comparison with COFlowNet, which is based on the flow matching (FM) objective (Bengio et al., 2021), we also adopt the FM objective (Equation (8)) in our experiments. Likewise, the following experiments are conducted under this same setting (results with additional objectives can be found in Appendix D.7). Moreover, we train GAIL using the rebalanced dataset described in Section 3.1 (rebalanced GAIL), which improves its performance relative to training on the original dataset. Accordingly, we evaluate it using the imitation policy learned via Algorithm 2.
@@
Each algorithm is evaluated using two standard metrics (Bengio et al., 2021; Pan et al., 2022; Zhang et al., 2023b; 2025b). The first is Empirical L1 Error, which measures the $\mathcal{L}_{1}$ distance between the empirical distribution $\pi(x)$ from model samples and the normalized reward distribution $p(x)=R(x)/Z$, computed as $\mathbb{E}[|\pi(x)-p(x)|]$. The second is Modes Found, defined as the number of distinct reward modes discovered during training.
We construct an Expert dataset of $1,500$ trajectories using a well-trained GFlowNet policy, simulating data that would otherwise be collected by a domain expert. It is important to clarify that we adopt GFlowNet-generated datasets not because our approach requires them, but because they provide a controlled and reproducible way to simulate realistic scenarios in which data arises from behavior policies of different optimality levels. This strategy is also common in the offline reinforcement learning community (Fu et al., 2020; Gulcehre et al., 2020). As demonstrated in the Section 4.2, TD-GFN also delivers compelling results on a real-world dataset. A more detailed discussion on this topic, as well as an analysis of trajectory structure’s value, is provided in Appendix E.
As shown in Figure 4(a), TD-GFN demonstrates superior sample efficiency. It identifies all $16$ modes with under $5,000$ state visits—a six-fold improvement over baselines—and converges to the lowest Empirical L1 Error more than twice as fast as the next best method, COFlowNet. A complete analysis of the other baselines can be found in Appendix D.8.
@@
Figure 4: Performance comparison on the $8^{4}$ Hypergrid task. The top row shows results using $1,500$ trajectories from the Expert and Mixed datasets for policy training, while the bottom row shows results using only $150$ trajectories. The solid line and shaded region represent the mean and standard deviation, respectively; the same convention applies to the following figures.
Robustness to Dataset Uncertainty.
@@
We first assess the robustness of TD-GFN under dataset uncertainty by introducing greater variability in the training data. In the Mixed setting, we augment the Expert dataset with an equal number of trajectories generated by a random policy similar to Hong et al. (2023); Cao et al. (2024), thereby introducing noise. In another setting, we significantly reduce the number of available trajectories—using only one-tenth of the original dataset—to simulate data scarcity as an additional source of uncertainty (Depeweg et al., 2018).
@@
Robustness to Behavior Policy Quality.
@@
Figure 5: Performance comparison on the $8^{4}$ Hypergrid task. Policies are trained using the Median and Bad datasets, each consisting of $1,500$ trajectories.
@@ 4.2 Biosequence Design
We next evaluate TD-GFN on a more challenging task—designing anti-microbial peptides (AMPs), i.e., short protein sequences—introduced in Jain et al. (2022). The experimental setup leverages two datasets curated from the DBAASP database (Pirtskhalava et al., 2021): one for training a reward proxy and another for an oracle that provides ground-truth labels.
Using the proxy reward model provided in Jain et al. (2022), we train a Proxy-GFN and compare its performance with TD-GFN, as well as with other baseline methods previously evaluated on the Hypergrid task (Section 4.1). All methods are trained on the same dataset used for proxy learning, which includes $3,219$ positive AMPs and $4,761$ non-AMPs.
Figure 6: Comparison of reward and diversity among the top $100$ sequences ranked by reward. The Diversity metric is defined in Appendix C.2. All results are reported over three random seeds.
@@ 4.3 Molecule Design
We further evaluate our method on the real-world molecular design benchmark introduced by Bengio et al. (2021). We use a training dataset that simulates historical expert data, consisting of $1,500$ trajectories generated by a moderately-trained GFlowNet policy (Behavior-GFN) and rewarded by the oracle provided in the original paper which was pre-trained on $300k$ molecules.
Table 1: Comparison of methods on the Molecule Design task. The Convergence column indicates the number of trajectories required for performance convergence. Results are reported as mean $\pm$ standard deviation over three random seeds, and bold numbers indicate the best performance in each column. The “Dataset” row shows the information of the $1,500$ samples in the training dataset for reference.
@@
Dataset $7.420\pm 0.088$ $6.968\pm 0.226$ $5.757\pm 0.635$ /
Behavior-GFN $7.534\pm 0.066$ $7.220\pm 0.165$ $6.504\pm 0.348$ /
Oracle-GFN $7.718\pm 0.014$ $7.408\pm 0.021$ $6.801\pm 0.023$ $44.141\times 10^{4}$
CQL $7.069$ $6.643$ $5.401$ 0.803 $\bm{\times 10^{4}}$
@@
TD-GFN (Ours) 7.733 $\pm$ 0.036 7.450 $\pm$ 0.037 6.810 $\pm$ 0.035 2.749 $\bm{\times 10^{4}}$
Following the architecture and methodology commonly adopted in prior GFlowNet research (Bengio et al., 2021; Malkin et al., 2022; Jang et al., 2023; Tiapkin et al., 2024), we train a proxy model on this dataset, which is then used to train a GFlowNet baseline referred to as Proxy-GFN. We also train an Oracle-GFN that uses the oracle model itself as the best-case proxy. We evaluate both the original (FM) COFlowNet and its quantile-augmented (QM) variant, proposed by Zhang et al. (2025b), which uses quantile matching (Zhang et al., 2023b) to improve candidate diversity in this task.
We report the average top-$k$ reward obtained from $5,000$ samples generated by policies trained with different algorithms, along with the number of training trajectories required for performance convergence, in Table 1. Due to the instability of offline RL methods under sparse rewards, results for CQL and IQL are reported using only their best-performing random seed.
As shown in the table, CQL converges fastest in terms of sample efficiency but struggles to generate a diverse set of high-reward candidates, resulting in substantially lower top-$k$ average rewards than GFlowNet-based and even imitation learning methods. Excluding CQL, TD-GFN achieves the best performance in discovering high-reward molecules with the fewest trajectories. Additional comparisons of training time and computational resource usage are provided in Appendix D.5.
Figure 7: Number of high-reward modes (reward >7>7 and >7.5>7.5) discovered during training in the Molecule Design task.
Notably, the powerful guidance provided at intermediate steps allows TD-GFN to match Oracle-GFN’s performance using $\mathbf{20}$ times fewer trajectories, while also decisively surpassing the weaker data-generating policy. This remarkable efficiency showcases its ability to transcend the limitation of training data and signals strong potential for iterative policy evolution.
To evaluate the diversity of molecules generated, we measure the number of distinct high-reward modes identified by each algorithm. Modes are clusters of high-reward molecules based on Tanimoto similarity (Bajusz et al., 2015). As shown in Figure 7, TD-GFN consistently discovers $1.5$ to $2$ times more high-reward modes than the strongest baselines, indicating superior exploration. Notably, despite incorporating DAG pruning, TD-GFN does not overfit to narrow regions of the search space. Instead, it guides the policy toward a broader set of high-reward regions, balancing targeted exploitation with diverse exploration.
@@ 5 Conclusion
We introduced Trajectory-Distilled GFlowNet (TD-GFN), a proxy-free training framework for learning GFlowNets from offline trajectory datasets without requiring out-of-distribution reward queries. Leveraging edge-level guidance learned via IRL, TD-GFN consistently steers the policy toward improved performance and training efficiency through DAG pruning and prioritized backward trajectory sampling. Extensive experiments demonstrate state-of-the-art performance and efficiency.
@@ Appendix A Algorithm Pseudocode
1:Input: Offline dataset $\mathcal{D}=\{\tau_{i}=(s_{0},\dots,s_{T_{i}},R(s_{T_{i}}))\}_{i=1}^{M}$; number of training iterations NN; batch size $B$; GFlowNet learning rate $\eta$.
@@
6:Prune the environment DAG $G=(V,E)$ based on the thresholding rule defined in Equation (6).
@@
9:Initialize the parameterized GFlowNet policy $\mathcal{P}_{F}$.
10:for $n=1$ to NN do
@@
16:     Sample $s_{pre}\sim\mathcal{P}_{B}(\cdot|s_{now})$, where $\mathcal{P}_{B}$ is defined in Equation (7) over $G^{\prime}$.
@@
19:   end while // Sample training trajectories via prioritized backward process.
@@
21:  end for
22:  Optimize $\mathcal{P}_{F}$ over $G^{\prime}$ with sampled trajectories $\mathcal{T}$. // Update policy.
@@
24:Output: Trained GFlowNet policy $\mathcal{P}_{F}$.
@@
6:   Sample a trajectory $\tau_{j}\in\mathcal{D}$ with probability $P(\tau_{j})\propto R(s_{T_{j}})$ .
@@
8:  end for // Sample from the rebalanced dataset.
@@
10:  Compute $\nabla_{\phi}L$ on $\overline{E}$; update $\phi\leftarrow\phi-\beta\nabla_{\phi}L$. // Update discriminator.
11:  Compute $\nabla_{\psi}L$ on $\overline{E}$; update $\psi\leftarrow\psi-\alpha\nabla_{\psi}L$ or apply a policy gradient method. // Update imitation policy.
@@
13:Extract the edge reward function $R_{E}$ according to Equation (5).
@@
Appendix B Related work